# Quantitative Evaluation of Multiple Instance Learning Reliability For WSIs Classification

### v2

@@ 3.3.1 CAMELYON16
The results for this dataset are presented in Tables 1 and 2. In Table 1, the ACMIL/2 model demonstrates moderate classification performance but high reliability. Conversely, DTFD achieves strong classification performance with lower reliability. Interestingly, the MAX-POOL model, while computationally efficient and yielding reasonable classification results, exhibits the lowest reliability. This finding reinforces our argument that model selection should not be based solely on classification metrics. To illustrate this point, we present a tumor slide that was correctly classified by MAX-POOL, along with its heatmap in Figure LABEL:fig:cam_max. The heatmap reveals that the model concentrates on normal patches of lymph node tissue rather than metastatic tumor patches, indicating the unreliability of its predictions.
The results in Table 2 show that additive models do not exhibit significant improvements in classification performance, but they do show slight enhancements in reliability metrics. Notably, the simple MEAN-POOL-INS model achieves the highest reliability values, although having the lowest classification performance. Figure LABEL:fig:cam_mean illustrates a slide and its corresponding heatmap, highlighting the model’s focus on tumor areas.
[…]
fig:cam_mean Refer to caption        Refer to caption
Figure 3: The test-40 slide with ground truth annotation (green) along with its heatmap generated by MEAN-POOL-INS.
@@
Table 3: The average reliability, classification, and computational metrics (± standard deviation) over five repetitions for CATCH.
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL 0.3015 ± 0.0185 0.5346 ± 0.0090 0.9103 ± 0.0071 0.987 ± 0.002 0.902 ± 0.015 94.4 M 791.3 K CLAM 0.3234 ± 0.0063 0.5383 ± 0.0036 0.9139 ± 0.0021 0.987 ± 0.002 0.881 ± 0.014 94.4 M 798.4 K DTFD 0.2007 ± 0.0599 0.5616 ± 0.0077 0.9141 ± 0.0054 0.995 ± 0.001 0.930 ± 0.011 126.1 M 1841.9 K ACMIL/4 0.2812 ± 0.0108 0.5546 ± 0.0083 0.9180 ± 0.0055 0.995 ± 0.001 0.926 ± 0.010 94.5 M 806.4 K MADMIL/2 0.2983 ± 0.0126 0.5500 ± 0.0042 0.9178 ± 0.0021 0.985 ± 0.001 0.904 ± 0.024 78.7 M 660.2 K MAX-POOL 0.0033 ± 0.0002 -0.0494 ± 0.0032 0.6540 ± 0.0041 0.998 ± 0.000 0.943 ± 0.015 62.9 M 528.3 K MEAN-POOL - - - 0.983 ± 0.001 0.809 ± 0.016 62.9 M 528.3 K
3.3.2 CATCH
Tables 3 and 4 present the results for CATCH dataset. Notably, the MAX-POOL model achieves the highest classification metrics but records the lowest reliability scores. Selecting the MAX-POOL model without considering its reliability could lead to unpredictable outcomes. As shown in Figure LABEL:fig:cat_max, the model primarily focuses on non-tumorous regions, further highlighting its lack of reliability.
Additionally, different models excel in various reliability metrics. Among them, the ACMIL/4 models (both plain and additive) strike a good balance between classification and reliability performance, although they are computationally demanding. In Figure LABEL:fig:cat_acmil, the heatmap produced by the plain ACMIL/4 model demonstrates that the model accurately focuses on regions of interest as annotated by pathologists.
Finally, the MEAN-POOL-INS model, while being the most computationally efficient, exhibits the lowest F1 score but achieves moderate reliability metrics
[…]
fig:cat_max Refer to caption        Refer to caption
Figure 4: A slide from CATCH with ground truth annotation (green) along with its heatmap generated by MAX-POOL.
@@ 3.3.2 CATCH
Table 4: The average reliability, classification, and computational metrics (± standard deviation) over five repetitions for the CATCH using additive models.
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL-ADD-ATT 0.2948 ± 0.0197 0.5266 ± 0.0030 0.9044 ± 0.0049 0.983 ± 0.003 0.896 ± 0.017 94.8 M 791.3 K ABMIL-ADD-PATCH 0.0589 ± 0.0068 0.3465 ± 0.0153 0.8253 ± 0.0073 CLAM-ADD-ATT 0.3077 ± 0.0218 0.5278 ± 0.0080 0.9014 ± 0.0050 0.983 ± 0.003 0.890 ± 0.016 94.8 M 798.5 K CLAM-ADD-PATCH 0.0672 ± 0.0023 0.3467 ± 0.0139 0.8262 ± 0.0056 DTFD-PATCH 0.0920 ± 0.0047 0.4187 ± 0.0085 0.8678 ± 0.0041 0.995 ± 0.001 0.930 ± 0.011 126.1 M 1841.9 K ACMIL/4-ADD-ATT 0.2280 ± 0.0796 0.5455 ± 0.0115 0.9086 ± 0.0066 0.993 ± 0.001 0.916 ± 0.014 94.9 M 806.4 K ACMIL/4-ADD-PATCH 0.1726 ± 0.0121 0.4438 ± 0.0169 0.8783 ± 0.0063 MAX-POOL-INS 0.0004 ± 0.0001 0.0043 ± 0.0068 0.6645 ± 0.0037 0.987 ± 0.005 0.885 ± 0.006 63.3 M 528.4 K MAX-POOL-INS-PATCH 0.2926 ± 0.0116 0.4720 ± 0.0581 0.8764 ± 0.0298 MEAN-POOL-INS 0.3194 ± 0.0034 0.4521 ± 0.0077 0.8525 ± 0.0039 0.983 ± 0.001 0.809 ± 0.016 63.3 M 528.4 K
[…]
fig:cat_acmil Refer to caption        Refer to caption
Figure 5: A slide from CATCH with ground truth annotation (green) along with its heatmap generated by ACMIL/4.
3.3.3 TCGA BRCA
The MADMIL/3 models exhibit the highest reliability performance, a high AUC classification metric, and low computational costs for this dataset, as shown in Table 5. Figure LABEL:fig:brca_madmil visually supports these numerical results. On the other hand, the DTFD model achieves the highest F1 score but demonstrates lower reliability, AUC, and computational efficiency. Additionally, we present a slide’s heatmap generated by the MEAN-POOL-INS model in Figure LABEL:fig:brca_mean-pool-ins, which aligns with the model’s moderate reliability, high classification AUC, and low computational cost reported in Table 6.
@@ 3.3.3 TCGA BRCA
Table 5: The average reliability, classification, and computational metrics (± standard deviation) over five repetitions for TCGA BRCA.
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL 0.1858 ± 0.0384 0.4358 ± 0.0360 0.8206 ± 0.0163 0.953 ± 0.008 0.853 ± 0.022 94.4 M 788.7 K CLAM 0.1779 ± 0.0138 0.4844 ± 0.0129 0.8388 ± 0.0088 0.951 ± 0.007 0.835 ± 0.011 94.4 M 790.7 K DTFD 0.1524 ± 0.0313 0.4748 ± 0.0104 0.8380 ± 0.0081 0.955 ± 0.006 0.869 ± 0.021 125.8 M 1052.6 K ACMIL/3 0.2211 ± 0.0119 0.4908 ± 0.0184 0.8498 ± 0.0113 0.961 ± 0.006 0.852 ± 0.018 94.5 M 792.3 K MADMIL/3 0.2238 ± 0.0155 0.4928 ± 0.0254 0.8530 ± 0.0164 0.966 ± 0.008 0.847 ± 0.040 73.5 M 614.8 K MAX-POOL 0.0038 ± 0.0003 -0.0639 ± 0.0045 0.6109 ± 0.0045 0.932 ± 0.003 0.796 ± 0.023 62.9 M 525.8 K MEAN-POOL - - - 0.977 ± 0.006 0.855 ± 0.017 62.9 M 525.8 K
[…]
fig:brca_madmil Refer to caption        Refer to caption
Figure 6: A slide from TCGA BRCA with ground truth annotation (green) along with corresponding heatmap generated by MADMIL/3.
@@
Table 6: The average reliability, classification, and computational metrics (± standard deviation) over five repetitions for TCGA BRCA using additive models.
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL-ADD-ATT 0.1907 ± 0.0256 0.4091 ± 0.1029 0.8116 ± 0.0418 0.910 ± 0.036 0.798 ± 0.035 94.5 M 788.7 K ABMIL-ADD-PATCH 0.0746 ± 0.0287 0.3265 ± 0.0264 0.8034 ± 0.0149 CLAM-ADD-ATT 0.1423 ± 0.0379 0.4770 ± 0.0242 0.8385 ± 0.0179 0.917 ± 0.011 0.786 ± 0.063 94.5 M 790.7 K CLAM-ADD-PATCH 0.0621 ± 0.0187 0.2863 ± 0.0377 0.7889 ± 0.0021 DTFD-PATCH 0.0667 ± 0.0085 0.2949 ± 0.0162 0.7942 ± 0.0058 0.955 ± 0.006 0.869 ± 0.021 125.8 M 1052.6 K ACMIL/3-ADD-ATT 0.1248 ± 0.0122 0.2254 ± 0.2965 0.7324 ± 0.1197 0.937 ± 0.010 0.843 ± 0.020 94.6 M 792.3 K ACMIL/3-ADD-PATCH 0.0587 ± 0.0056 0.1125 ± 0.0750 0.7026 ± 0.0391 MAX-POOL-INS 0.0002 ± 0.0001 0.0104 ± 0.0070 0.6132 ± 0.0043 0.894 ± 0.040 0.725 ± 0.057 63.0 M 525.8 K MAX-POOL-INS-PATCH 0.2364 ± 0.0427 0.4495 ± 0.0403 0.8392 ± 0.0165 MEAN-POOL-INS 0.2672 ± 0.0027 0.3037 ± 0.0107 0.7873 ± 0.0054 0.976 ± 0.006 0.850 ± 0.018 63.0 M 525.8 K
[…]
fig:brca_mean-pool-ins Refer to caption        Refer to caption
@@
Figure 7: A slide from TCGA BRCA with ground truth annotation (green) along with corresponding heatmap generated by MEAN-POOL-INS.
@@
Model Reliability Classification Computation MI Spearman’s AUPRC AUC F1 FLOPs Model Size ABMIL 0.2013 0.4195 0.7719 0.929 0.866 94.4 M 789.5 K CLAM 0.2063 0.4384 0.7805 0.932 0.855 94.4 M 793.2 K DTFD 0.1602 0.4434 0.7859 0.947 0.882 125.9 M 1315.7 K ACMIL 0.2211 0.4602 0.7844 0.926 0.862 94.4 M 796.5 K MADMIL 0.2183 0.4495 0.7929 0.937 0.866 76.9 M 644.2 K MAX-POOL 0.0029 -0.0377 0.4967 0.909 0.844 62.9 M 526.6 K MEAN-POOL - - - 0.847 0.751 62.9 M 526.6 K
3.4 Discussion
Based on the overall mean of the average results across the three datasets, as presented in Tables 7, 8, and Figure LABEL:fig:bar_plot, several key conclusions can be drawn. First, while the simple MAX-POOL network achieves high classification metrics and low computational cost, it demonstrates poor reliability values, making it unsuitable for real-world applications. In contrast, the MEAN-POOL-INS model shows high reliability, acceptable classification performance, and low computational cost, making it a more favorable choice.
@@ 3.4 Discussion
The multihead networks, ACMIL and MADMIL, offer both high reliability and classification performance, but at a higher computational expense, particularly ACMIL. Therefore, they are suitable to have a fair balance among different metrics. Meanwhile, the DTFD model delivers the highest classification performance, though this comes at the cost of increased computational demands and reduced reliability. Therefore, when classification performance is a priority, and computational cost and reliability are secondary concerns, DTFD may be a suitable choice. However, in scenarios where computational efficiency and reliability are more critical than classification performance, the MEAN-POOL-INS model is recommended.
@@
Regarding the additive models, while they do not exhibit a significant improvement in classification metrics, they do show a subtle increase in reliability scores, albeit with slightly higher computational cost. Although additive models tend to be more interpretable, they are not necessarily more reliable.
At the end, we urge researchers to include computational cost and reliability metrics alongside classification results when evaluating and developing new models. Our findings demonstrate that classification metrics alone do not fully reflect the diverse capabilities of models, as evidenced by the results in this study.
3.5 Conclusion

### v3

@@ Results
The results for this dataset are presented in Tables  1 and 2. In Table  1, the ACMIL/2 model demonstrates moderate classification performance but high reliability. Conversely, DTFD achieves strong classification performance with lower reliability. Interestingly, the MAX-POOL model, while computationally efficient and yielding reasonable classification results, exhibits the lowest reliability. This finding reinforces our argument that model selection should not be based solely on classification metrics. To illustrate this point, we present a tumor slide that was correctly classified by MAX-POOL, along with its heatmap in Fig 2. The heatmap reveals that the model concentrates on normal patches of lymph node tissue rather than metastatic tumor patches, indicating the unreliability of its predictions.
@@
Table 1: Average reliability, classification, and computational metrics (± standard deviation) over five repetitions for the CAMELYON16 dataset.
Model Reliability Classification Computation
MI Spearman’s AUPRC AUC F1 FLOPs Size
ABMIL 0.12 ± 0.00 0.29 ± 0.00 0.59 ± 0.01 0.85 ± 0.01 0.84 ± 0.01 94 M 789 K
CLAM 0.12 ± 0.01 0.29 ± 0.02 0.59 ± 0.02 0.86 ± 0.01 0.85 ± 0.01 94 M 791 K
DTFD 0.13 ± 0.01 0.29 ± 0.01 0.61 ± 0.01 0.89 ± 0.01 0.85 ± 0.02 126 M 1053 K
ACMIL/2 0.16 ± 0.02 0.34 ± 0.02 0.59 ± 0.05 0.82 ± 0.03 0.81 ± 0.01 94 M 791 K
MADMIL/2 0.13 ± 0.00 0.31 ± 0.00 0.61 ± 0.00 0.86 ± 0.01 0.85 ± 0.01 79 M 658 K
MAX-POOL 0.00 ± 0.00 -0.00 ± 0.00 0.23 ± 0.01 0.80 ± 0.01 0.79 ± 0.01 63 M 526 K
MEAN-POOLa 0.58 ± 0.02 0.59 ± 0.02 63 M 526 K
FLOPs are computed with 120 instances per bag. The highest value in each column is bolded using the exact value, while the tables report numbers rounded to two decimals for readability.
a Reliability metrics are not computed for this model, as it does not provide patch scores.
Fig 2: (I) The test-30 slide with ground truth annotations (green) overlaid on the tissue section. (II) Corresponding heatmap generated by MAX-POOL, showing predicted patch scores distribution from low (blue) to high (red). The annotation and heatmap are spatially aligned for comparison.
@@
Table 2: Average reliability, classification, and computational metrics (± standard deviation) over five repetitions for the CAMELYON16 dataset using additive models.
Model Reliability Classification Computation
MI Spearman’s AUPRC AUC F1 FLOPs Size
ABMIL-ADD-ATT 0.13 ± 0.00 0.32 ± 0.01 0.62 ± 0.01 0.86 ± 0.01 0.83 ± 0.02 95 M 789 K
ABMIL-ADD-PATCH 0.04 ± 0.00 0.24 ± 0.01 0.48 ± 0.01
CLAM-ADD-ATT 0.12 ± 0.01 0.29 ± 0.01 0.60 ± 0.02 0.87 ± 0.01 0.83 ± 0.02 95 M 791 K
CLAM-ADD-PATCH 0.03 ± 0.00 0.24 ± 0.05 0.48 ± 0.02
DTFD-PATCHa 0.04 ± 0.00 0.29 ± 0.02 0.53 ± 0.01 0.89 ± 0.01 0.85 ± 0.02 126 M 1053 K
ACMIL/2-ADD-ATT 0.17 ± 0.01 0.35 ± 0.01 0.62 ± 0.02 0.79 ± 0.02 0.81 ± 0.01 95 M 791 K
ACMIL/2-ADD-PATCH 0.03 ± 0.00 -0.03 ± 0.01 0.29 ± 0.01
MAX-POOL-INS 0.00 ± 0.00 0.05 ± 0.03 0.20 ± 0.02 0.77 ± 0.14 0.76 ± 0.13 63 M 526 K
MAX-POOL-INS-PATCH 0.15 ± 0.06 0.29 ± 0.15 0.55 ± 0.20
MEAN-POOL-INS 0.31 ± 0.04 0.53 ± 0.07 0.78 ± 0.08 0.58 ± 0.02 0.59 ± 0.02 63 M 526 K
a We use the provided patch probabilities based on GRAD-CAM.
The results in Table 2 show that additive models do not exhibit significant improvements in classification performance, but they do show slight enhancements in reliability metrics. Notably, the simple MEAN-POOL-INS model achieves the highest reliability values, although having the lowest classification performance. Fig 3 illustrates a slide and its corresponding heatmap, highlighting the model’s focus on tumor areas.
Fig 3: (I) The test-40 slide with ground truth annotations (green) overlaid on the tissue section. (II) Corresponding heatmap generated by MEAN-POOL-INS, showing predicted patch scores distribution from low (blue) to high (red). The annotation and heatmap are spatially aligned for comparison.
CATCH
Tables 3 and 4 present the results for CATCH dataset. Notably, the MAX-POOL model achieves the highest classification metrics but records the lowest reliability scores. Selecting the MAX-POOL model without considering its reliability could lead to unpredictable outcomes. As shown in Fig 4, the model primarily focuses on non-tumorous regions, further highlighting its lack of reliability.
@@
Table 3: Average reliability, classification, and computational metrics (± standard deviation) over five repetitions for CATCH.
Model Reliability Classification Computation
MI Spearman’s AUPRC AUC F1 FLOPs Size
ABMIL 0.30 ± 0.02 0.53 ± 0.01 0.91 ± 0.01 0.99 ± 0.00 0.90 ± 0.02 94 M 791 K
CLAM 0.32 ± 0.01 0.54 ± 0.00 0.91 ± 0.00 0.99 ± 0.00 0.88 ± 0.01 94 M 798 K
DTFD 0.20 ± 0.06 0.56 ± 0.01 0.91 ± 0.01 1 ± 0.00 0.93 ± 0.01 126 M 1842 K
ACMIL/4 0.28 ± 0.01 0.55 ± 0.01 0.92 ± 0.01 1 ± 0.00 0.93 ± 0.01 95 M 807 K
MADMIL/2 0.30 ± 0.01 0.55 ± 0.00 0.92 ± 0.00 0.99 ± 0.00 0.90 ± 0.02 79 M 660 K
MAX-POOL 0.00 ± 0.00 -0.05 ± 0.00 0.65 ± 0.00 1 ± 0.000 0.94 ± 0.02 63 M 528 K
MEAN-POOL - - - 0.98 ± 0.00 0.81 ± 0.02 63 M 528 K
@@
Table 4: Average reliability, classification, and computational metrics (± standard deviation) over five repetitions for the CATCH using additive models.
Model Reliability Classification Computation
MI Spearman’s AUPRC AUC F1 FLOPs Size
ABMIL-ADD-ATT 0.29 ± 0.02 0.53 ± 0.00 0.90 ± 0.00 0.98 ± 0.00 0.90 ± 0.02 95 M 791 K
ABMIL-ADD-PATCH 0.06 ± 0.01 0.35 ± 0.02 0.83 ± 0.01
CLAM-ADD-ATT 0.31 ± 0.02 0.53 ± 0.01 0.90 ± 0.01 0.98 ± 0.00 0.89 ± 0.02 95 M 799 K
CLAM-ADD-PATCH 0.07 ± 0.00 0.35 ± 0.01 0.83 ± 0.01
DTFD-PATCH 0.09 ± 0.00 0.42 ± 0.01 0.87 ± 0.00 1 ± 0.00 0.93 ± 0.01 126 M 1842 K
ACMIL/4-ADD-ATT 0.23 ± 0.08 0.55 ± 0.01 0.91 ± 0.01 0.99 ± 0.00 0.92 ± 0.01 95 M 806 K
ACMIL/4-ADD-PATCH 0.17 ± 0.01 0.44 ± 0.02 0.88 ± 0.01
MAX-POOL-INS 0.00 ± 0.00 0.00 ± 0.01 0.66 ± 0.00 0.99 ± 0.01 0.89 ± 0.01 63 M 528 K
MAX-POOL-INS-PATCH 0.29 ± 0.01 0.47 ± 0.06 0.88 ± 0.03
MEAN-POOL-INS 0.32 ± 0.00 0.45 ± 0.01 0.85 ± 0.00 0.98 ± 0.00 0.81 ± 0.02 63 M 528 K
Fig 4: (I) A slide from CATCH with ground truth annotations (green) overlaid on the tissue section. (II) Corresponding heatmap generated by MAX-POOL, showing predicted patch scores distribution from low (blue) to high (red). The annotation and heatmap are spatially aligned for comparison.
Additionally, different models excel in various reliability metrics. Among them, the ACMIL/4 models (both plain and additive) strike a good balance between classification and reliability performance, although they are computationally demanding. In Fig 5, the heatmap produced by the plain ACMIL/4 model demonstrates that the model accurately focuses on regions of interest as annotated by pathologists.
Fig 5: (I) A slide from CATCH with ground truth annotations (green) overlaid on the tissue section. (II) Corresponding heatmap generated by ACMIL/4, showing predicted patch scores distribution from low (blue) to high (red). The annotation and heatmap are spatially aligned for comparison.
Finally, the MEAN-POOL-INS model, while being the most computationally efficient, exhibits the lowest F1 score but achieves moderate reliability metrics.
TCGA BRCA
The MADMIL/3 models exhibit the highest reliability performance, a high AUC classification metric, and low computational costs for this dataset, as shown in Table 5.
Table 5: Average reliability, classification, and computational metrics (± standard deviation) over five repetitions for TCGA BRCA.
Model Reliability Classification Computation
MI Spearman’s AUPRC AUC F1 FLOPs Size
ABMIL 0.19 ± 0.04 0.44 ± 0.04 0.82 ± 0.02 0.95 ± 0.01 0.85 ± 0.02 94 M 789 K
CLAM 0.18 ± 0.01 0.48 ± 0.01 0.84 ± 0.01 0.95 ± 0.01 0.84 ± 0.01 94 M 791 K
DTFD 0.15 ± 0.03 0.47 ± 0.01 0.84 ± 0.01 0.96 ± 0.01 0.87 ± 0.02 126 M 1053 K
ACMIL/3 0.22 ± 0.01 0.49 ± 0.02 0.85 ± 0.01 0.96 ± 0.01 0.85 ± 0.02 95 M 792 K
MADMIL/3 0.22 ± 0.02 0.49 ± 0.03 0.85 ± 0.02 0.97 ± 0.01 0.85 ± 0.04 74 M 615 K
MAX-POOL 0.00 ± 0.00 -0.06 ± 0.00 0.61 ± 0.00 0.93 ± 0.00 0.80 ± 0.02 63 M 526 K
MEAN-POOL - - - 0.98 ± 0.01 0.86 ± 0.02 63 M 526 K
Fig 6 visually supports these numerical results. On the other hand, the DTFD model achieves the highest F1 score but demonstrates lower reliability, AUC, and computational efficiency. Additionally, we present a slide’s heatmap generated by the MEAN-POOL-INS model in Fig 7, which aligns with the model’s moderate reliability, high classification AUC, and low computational cost reported in Table 6.
Fig 6: (I) A slide from TCGA BRCA with ground truth annotations (green) overlaid on the tissue section. (II) Corresponding heatmap generated by MADMIL/3, showing predicted patch scores distribution from low (blue) to high (red). The annotation and heatmap are spatially aligned for comparison.
@@
Fig 7: (I) A slide from TCGA BRCA with ground truth annotations (green) overlaid on the tissue section. (II) Corresponding heatmap generated by MEAN-POOL-INS, showing predicted patch scores distribution from low (blue) to high (red). The annotation and heatmap are spatially aligned for comparison.
Table 6: Average reliability, classification, and computational metrics (± standard deviation) over five repetitions for TCGA BRCA using additive models.
Model Reliability Classification Computation
MI Spearman’s AUPRC AUC F1 FLOPs Size
ABMIL-ADD-ATT 0.19 ± 0.03 0.41 ± 0.10 0.81 ± 0.04 0.91 ± 0.04 0.80 ± 0.04 95 M 789 K
ABMIL-ADD-PATCH 0.07 ± 0.03 0.33 ± 0.03 0.80 ± 0.01
CLAM-ADD-ATT 0.14 ± 0.04 0.48 ± 0.02 0.84 ± 0.02 0.92 ± 0.01 0.79 ± 0.06 95 M 791 K
CLAM-ADD-PATCH 0.06 ± 0.02 0.29 ± 0.04 0.79 ± 0.00
DTFD-PATCH 0.07 ± 0.01 0.29 ± 0.02 0.79 ± 0.01 0.96 ± 0.01 0.87 ± 0.02 126 M 1053 K
ACMIL/3-ADD-ATT 0.12 ± 0.01 0.23 ± 0.30 0.73 ± 0.12 0.94 ± 0.01 0.84 ± 0.02 95 M 792 K
ACMIL/3-ADD-PATCH 0.06 ± 0.01 0.11 ± 0.08 0.70 ± 0.04
MAX-POOL-INS 0.00 ± 0.00 0.01 ± 0.01 0.61 ± 0.00 0.89 ± 0.04 0.73 ± 0.06 63 M 526 K
MAX-POOL-INS-PATCH 0.24 ± 0.04 0.45 ± 0.04 0.84 ± 0.02
MEAN-POOL-INS 0.27 ± 0.00 0.30 ± 0.01 0.79 ± 0.01 0.98 ± 0.01 0.85 ± 0.02 63 M 526 K
Discussion
Based on the overall mean of the average results across the three datasets, as presented in Tables 78, and Fig 8, several key conclusions can be drawn. Repeated-measures ANOVA revealed significant differences across models for all evaluated metrics. For each model the three datasets were treated as repeated measures. Specifically, significant effects of model were observed on MI ($F(5,20)=178.47$, $p<0.001$), Spearman correlation ($F(5,20)=12.34$, $p=0.0001$), AUPRC ($F(5,20)=9.12$, $p=0.0005$), AUC ($F(6,24)=15.21$, $p<0.001$), and F1 score ($F(6,24)=8.45$, $p=0.0012$), indicating that model performances differed significantly across datasets. The learning curves for all methods on the TCGA BRCA dataset are provided in the Supporting information. These curves illustrate the convergence dynamics and stability across folds, supporting the reported performance metrics.
The model comparisons highlight important trade-offs. First, while the simple MAX-POOL network achieves high classification metrics and low computational cost, it demonstrates poor reliability values. This inconsistency may compromise clinical trust, suggesting that despite its efficiency, MAX-POOL is unsuitable for deployment in real-world applications where reliability is critical. In contrast, the MEAN-POOL-INS model shows high reliability, comparatively good classification performance, and low computational cost, making it a more favorable choice for scenarios where consistent predictions and resource constraints are both important.
@@ Discussion
Model Reliability Classification Computation
MI Spearman’s AUPRC AUC F1 FLOPs Size
ABMIL 0.20 0.42 0.77 0.93 0.87 94 M 790 K
CLAM 0.21 0.44 0.78 0.93 0.86 94 M 793 K
DTFD 0.16 0.44 0.79 0.95 0.88 126 M 1316 K
ACMIL 0.22 0.46 0.78 0.93 0.86 94 M 797 K
MADMIL 0.22 0.45 0.79 0.94 0.87 77 M 644 K
MAX-POOL 0.00 -0.04 0.50 0.91 0.84 63 M 527 K
MEAN-POOL - - - 0.85 0.75 63 M 527 K
Model Reliability Classification Computation
MI Spearman’s AUPRC AUC F1 FLOPs Size
ABMIL-ADD-ATT 0.21 0.42 0.78 0.92 0.84 95 M 790 K
ABMIL-ADD-PATCH 0.06 0.30 0.70
CLAM-ADD-ATT 0.19 0.43 0.78 0.92 0.83 95 M 793 K
CLAM-ADD-PATCH 0.05 0.29 0.70
DTFD-PATCH 0.07 0.33 0.73 0.95 0.88 126 M 1316 K
ACMIL-ADD-ATT 0.18 0.37 0.75 0.91 0.86 95 M 797 K
ACMIL-ADD-PATCH 0.09 0.17 0.62
MAX-POOL-INS 0.00 0.02 0.49 0.88 0.79 63 M 527 K
MAX-POOL-INS-PATCH 0.23 0.40 0.76
MEAN-POOL-INS 0.30 0.43 0.81 0.85 0.75 63 M 527 K
Fig 8: Bar plots comparing the models with different metrics.
@@
The multihead networks, ACMIL and MADMIL, offer both high reliability and classification performance, but at a higher computational expense, highlighting their potential for research or clinical applications where both accuracy and consistency are essential. Their higher computational cost, particularly for ACMIL, indicates a trade-off that must be considered in resource-limited settings. Meanwhile, the DTFD model delivers the highest classification performance, though this comes at the cost of increased computational demands and reduced reliability. Therefore, when classification performance is a priority, and computational cost and reliability are secondary concerns, DTFD may be a suitable choice.
@@
Regarding the additive models, while they do not exhibit a significant improvement in classification metrics, they do show a subtle increase in reliability scores, albeit with slightly higher computational cost. While they offer enhanced interpretability, their limited benefits in reliability and accuracy indicate that they may be most valuable in research contexts emphasizing model explainability rather than operational performance.
To better understand the contribution of each reliability metric, we examined the model performance by selectively excluding MI, Spearman correlation, or AUPRC from the overall model ranking (Table 9). The results show that excluding MI substantially reduces the ability to differentiate models with consistent predictions, as several top models become tied, highlighting MI’s critical role in capturing prediction consistency. Excluding AUPRC also affects the rankings, emphasizing its importance in assessing realiability. Similarly, removing Spearman correlation impacts the overall ranking, causing several models to tie at lower positions. These findings indicate that all three metrics contribute meaningfully to assessing model reliability, and a multi-metric framework is necessary to capture different aspects of predictive behavior effectively.
Table 9: Analysis of reliability metrics showing the effect of excluding each metric on model ranking. Rankings are obtained by summing the scores of the selected reliability metrics (with equal weight) across CAMELYON16, CATCH, and TCGA BRCA datasets, and ranking models based on the aggregated score.
Metric Excluded Model Ranking (Reliability)
None (All metrics) MEAN-POOL-INS >> ACMIL >> MADMIL >> CLAM >> ABMIL-ADD-ATT >> CLAM-ADD-ATT
MI excluded MEAN-POOL-INS == ACMIL == MADMIL >> DTFD >> CLAM >> CLAM-ADD-ATT
Spearman’s excluded MEAN-POOL-INS >> MADMIL >> ACMIL >> CLAM == ABMIL-ADD-ATT == MAX-POOL-INS-PATCH
AUPRC excluded MEAN-POOL-INS >> ACMIL >> MADMIL >> CLAM >> ABMIL-ADD-ATT == MAX-POOL-INS-PATCH
Overall, these findings underscore that classification metrics alone do not capture the full utility of a model. Incorporating reliability and computational cost alongside accuracy provides a more comprehensive assessment, guiding model selection. While standard metrics such as AUC, accuracy, and F1 capture overall classification performance, they can fail in certain scenarios, such as rare positive instances, highly imbalanced classes, or heterogeneous MIL bags. In our experiments, the MAX-POOL network achieves high AUC and F1 scores but demonstrates low reliability, highlighting that strong classification metrics alone do not guarantee trustworthy predictions. Incorporating reliability metrics alongside standard metrics allows identification of models that not only perform well on average but also make consistent and robust predictions, providing a more comprehensive assessment of model behavior and guiding informed model selection. By explicitly considering these factors, researchers can design MIL architectures that balance predictive performance, computational efficiency, and deployability, ultimately advancing the development of models suited for real-world WSI analysis.
Conclusion